There are 150 students in a class. The distribution if their marks in a mathematics test are as follows



Class frequency



0-9 3



10-19 10



20-29 17



30-39 x



40-49 35



50-59 y



60-69 18



70-79 10



80-89 5



90-99 2



Required



i) The value of x given that the median mark is 44.357 (2marks)



ii) The value of y given that the modal is 43.0 (2marks)



iii) Draw an ogive of the data in (a) above (3 marks)

1
Expert's answer
2022-03-17T07:36:21-0400



i. median "=44.357=L+(\frac{(\frac{N}{2}-pcf)}{f})i"

"=40+(\frac{(75-(30-x)}{35})10"

"=39.5+(\frac{(45-x)}{35})10"

"4.357=(\frac{(45-x)}{35})10"

"0.4357=\frac{(45-x)}{35}"

"15.2495=45-x"

"x=29.75"

ii. Mode "=43.0=40+(\frac{7}{(7+35-y)})10"

"40+(\frac{7}{(42-y)})10"

"3=(\frac{7}{(42-y)})10"

"0.3=\frac{7}{(42-y)}"

"42-y=\frac{7}{0.3}=23.33"

"y=42-2.33=18.67"

iii.